OpenStudy (landon34):

Haley is at the store buying some supplies for an art project. She decides to buy some colored pencils for $3.95 and a drawing tablet. The total cost of the supplies before sales tax is $6.94. Which equation could not be used to find the cost of the drawing tablet? $3.95 + x = $6.94 $6.94 + x = $3.95 x = $6.94 - $3.95 $6.94 - x = $3.95

OpenStudy (landon34):

could you help me?

OpenStudy (anonymous):

The colored pencils (3.95) + the drawing tablet (x) has to be equal to 6.94. What equation is this?

OpenStudy (landon34):

3.95+x=6.94?

OpenStudy (anonymous):

Correct, that one CAN be used, now you can rewrite that equation to get 2 other equations. This leaves you with 1 there CANT be used.

OpenStudy (anonymous):

Try and rewrite that equation into something else

OpenStudy (landon34):

6.94+x=3.95 would work, right?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

That is the one there doesnt work.

OpenStudy (landon34):

so that is the answer then. thanks.

OpenStudy (landon34):

:D

OpenStudy (anonymous):

(which is the answer to this question). The 2 others works, we have:\[3.95+x=6.94\] Now we can subtract with x on both sides: \[3.95=6.95-x\] But we could also subtract with 3.95 \[x=6.94-3.95\]

OpenStudy (anonymous):

So yea, that is the answer

OpenStudy (landon34):

ok thanks.